# Quantum-annealing-inspired algorithms for multijet clustering

### v1

@@ 5 Conclusion and Outlook
In our study, jet clustering is formulated as a QUBO problem and three QAIAs are adopted to pursue global reconstruction, which is a new approach compared to the traditional iterative reconstruction, implemented in FastJet, for example. The distance defined in the QUBO design has a significant impact on the reconstruction performance, particularly when the jet multiplicity is beyond two. It is found that the angle-based approach only provides reasonable performance for dijet events and alternative distances such as $ee$-$k_{t}$ are mandatory for higher jet multiplicities. For such multijet cases, it becomes exceedingly difficult even to approximately search for the “quasi”-ground state, ending up in a local minimum, which has an order-of-magnitude higher energy than the optimal states (Figure 5 (c,e)). Because of such a challenge, previous studies saw fairly degraded performance in multijet reconstruction. However, this study shows that a powerful QUBO problem solver, namely bSB among the three QAIAs, can find optimal solutions and demonstrates its outstanding capability to solve combinatorial optimization problems, despite the high QUBO connectivities. It is promising to see that this global jet reconstruction with bSB can bring in improvement to the invariant mass reconstruction in multijet events.
Being a “quantum-inspired” algorithm, bSB runs on classical computers and is suitable for parallel processing and usage of cutting-edge computing resources such as GPUs and FPGAs. It is important to note that we can flexibly balance speed and reconstruction precision in QAIAs. Namely, if precise energy resolution is not required, for low-level triggers, for example, we can run the reconstruction much faster. Thus, with the applicability to run on FPGAs, bSB may particularly provide an important option to be considered for triggers, during the $Z$-pole data taking at the CEPC, an “exa-byte”-level data taking comparable to the intensive HL-LHC conditions. The running time tends to be longer for high jet multiplicity events, and further investigations on the speed-up are ongoing.
@@
Acknowledgements
HO would like to thank Gang Li, Shudong Wang and Xu Gao for feedback regarding the CEPC 4th detector concept in Delphes. HO is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12075060 and the NSFC Basic Science Center Program for Joint Research on High Energy Frontier Particle Physics under Grant No. 12188102.
Appendix A Simulated bifurcation
SB emulates quantum bifurcation phenomena within classical nonlinear Hamiltonian systems. The Hamiltonian of an Ising model with $N$ spin $\{x_{i}\}_{i=1}^{N}\in\{-1,1\}^{N}$ is defined as
$H(x_{i})=\frac{1}{2}\sum_{ij}^{N}J_{ij}x_{i}x_{j}+\sum_{i}^{N}h_{i}x_{i},$ (9)
which is equivalent to Eq. (1) and they are related by:
$x_{i}$ $\longleftrightarrow 2s_{i}-1$ (10)
$J_{ij}$ $\longleftrightarrow\frac{Q_{ij}}{2}$ (11)
$h_{i}$ $\longleftrightarrow\frac{\sum_{j}Q_{ij}}{2}$ (12)
The first version of SB algorithms proposed, adiabatic SB (aSB) [42], is formulated as:
$\centering\dot{x}_{i}\@add@centering$ $=$ $\frac{\partial H_{\textrm{aSB}}}{\partial y_{i}}=a_{0}y_{i}$ (13)
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{aSB}}}{\partial x_{i}},$ (14)
$=$ $-\left[x_{i}^{2}+a_{0}-a(t)\right]x_{i}+c_{0}h_{i}+c_{0}\sum_{j=1}^{N}J_{ij}x_{j},$
$H_{\textrm{aSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (15)
$V_{\textrm{aSB}}$ $=$ $\sum_{i=1}^{N}\left(\frac{x_{i}^{4}}{4}+\frac{a_{0}-a(t)}{2}x_{i}^{2}\right)$ (16)
$-c_{0}\sum_{i=1}^{N}h_{i}x_{i}-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}x_{j},$
where $x_{i}$ and $y_{i}$ are respectively the position and momentum of the $i$-th Kerr-nonlinear oscillator corresponding to the $i$-th spin, $a_{0}$ and $c_{0}$ are positive constants (the detuning frequency for the former and the coupling strength for the latter), $a(t)$ is a time-dependent pumping amplitude that monotonically increases from zero to $a_{0}$, $H_{\textrm{aSB}}$ is the Hamiltonian, and $V_{\textrm{aSB}}$ is the potential energy in aSB.
The ballistic SB (bSB) algorithm [43] introduces complete inelastic scattering walls at $|x_{i}|>1$; i.e. $x_{i}$ is forced to be $x_{i}=\pm 1$ and $y_{i}=0$ when the particle passes the threshold:
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{bSB}}}{\partial x_{i}},$ (18)
$=$ $-\left[a_{0}-a(t)\right]x_{i}+c_{0}h_{i}+c_{0}\sum_{j=1}^{N}J_{ij}x_{j},$
$H_{\textrm{bSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (19)
$V_{\textrm{bSB}}$ $=$ $\frac{a_{0}-a(t)}{2}\sum_{i=1}^{N}x_{i}^{2}-c_{0}\sum_{i=1}^{N}h_{i}x_{i}-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}x_{j}\quad$ (20)
$(\textrm{when}~{}|x_{i}|\leq 1~{}\textrm{for all}~{}x_{i}),$
$=$ $\infty\quad(\textrm{otherwise}).$
The quartic term is removed to avoid divergence, as the inelastic wall already plays similar roles to the nonlinear potential walls.
An alternative approach, discrete SB (dSB) [43], introduces discretization of the position, replacing $x_{i}$ with $\textrm{sgn}(x_{i})$:
$\centering\dot{x}_{i}\@add@centering$ $=$ $\frac{\partial H_{\textrm{dSB}}}{\partial y_{i}}=a_{0}y_{i}$ (21)
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{dSB}}}{\partial x_{i}},$ (22)
$=$ $-\left[a_{0}-a(t)\right]x_{i}+c_{0}h_{i}$
$+c_{0}\sum_{j=1}^{N}J_{ij}~{}\textrm{sgn}(x_{j}),$
$H_{\textrm{dSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (23)
$V_{\textrm{dSB}}$ $=$ $\frac{a_{0}-a(t)}{2}\sum_{i=1}^{N}x_{i}^{2}-c_{0}\sum_{i=1}^{N}h_{i}x_{i}$ (24)
$-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}~{}\textrm{sgn}(x_{j})\quad$
$(\textrm{when}~{}|x_{i}|\leq 1~{}\textrm{for all}~{}x_{i})$
$=$ $\infty\quad(\textrm{otherwise}),$
where the abovementioned partial discretization is introduced by replacing some $x_{i}$ with $\textrm{sgn}(x_{i})$.
Appendix B Ising energy predictions and evolution against running time
Figure 5 demonstrates the minimum Ising energies predicted by the three QAIAs for the three physics processes. bSB can achieve significantly improved predication of the minimum energies especially for high jet multiplicity events and QUBOs using the $ee$-$k_{t}$ distance. Figure 6 shows the evolution of Ising energies from the three QAIAs in the corresponding events.
Figure 5: Minimum Ising energies by the three QAIAs for a $Z$ boson (a,b), $ZH$ (c,d) and $t\bar{t}$ events (e,f) with the $ee$-$k_{t}$-based (a,c,e) or angle-based (b,d,f) distance considered in the QUBO formulation.
Figure 6: Evolution of Ising energies against the running time from the three QAIAs presented for (a) $Z$ boson, (b) $ZH$, and (c) $t\bar{t}$ events. The solid lines are the average of multiple shots (100 for $Z$ and 50 for $ZH$ and $t\bar{t}$) and the envelopes represent the standard deviation from the mean of those shots.
Appendix C Event displays
Event displays for the three physics processes are presented in Figures 7-9 reconstructed with the benchmark FastJet $ee$-$k_{t}$ algorithm or the bSB, dSB and D-Wave Neal using the $ee$-$k_{t}$ or the angled-based QUBOs.

### v2

@@ 5 Conclusion and Outlook
In our study, jet clustering is formulated as a QUBO problem. Three QAIAs are adopted to pursue global reconstruction, which is a new approach compared to the traditional iterative reconstruction, implemented in FastJet. The distance defined in the QUBO design significantly impacts the reconstruction performance, particularly when the jet multiplicity is beyond two. We find that the angle-based approach only provides reasonable performance for dijet events, and alternative distances such as $ee$-$k_{t}$ are mandatory for higher jet multiplicities. For such multijet cases, it becomes exceedingly difficult even to approximately search for the “quasi”-ground state, ending up in a local minimum, which has an order-of-magnitude higher energy than the optimal states (Figure 1(b)1(c)). Because of such a challenge, previous studies saw fairly degraded performance in multijet reconstruction. However, this study shows that a powerful QUBO problem solver, bSB among the three QAIAs, can find optimal solutions. It demonstrates its outstanding capability to solve combinatorial optimization problems despite the high QUBO connectivities. It is promising that this global jet reconstruction with bSB can improve the invariant mass reconstruction in multijet events.
As a “quantum-inspired” algorithm, bSB runs on classical computers and is suitable for parallel processing and using cutting-edge computing resources such as GPUs and FPGAs. It is important to note that we can flexibly balance speed and reconstruction precision in QAIAs. Namely, if precise energy resolution is not required for low-level triggers, for example, we can run the reconstruction much faster. Thus, with the applicability to run on FPGAs, bSB may particularly provide an important option to be considered for triggers during the $Z$-pole data taking at the CEPC, an “exabyte”-level data taking comparable to the intensive HL-LHC conditions. The running time tends to be longer for high jet multiplicity events, and further investigations on the speed-up are ongoing.
@@
Acknowledgments
HO would like to thank Gang Li, Shudong Wang, and Xu Gao for their feedback regarding the CEPC 4th detector concept in Delphes. HO is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12075060 and the NSFC Basic Science Center Program for Joint Research on High Energy Frontier Particle Physics under Grant No. 12188102.
Appendix A Event displays
Figures 6-8 present event displays for the three physics processes reconstructed with the benchmark FastJet $ee$-$k_{t}$ algorithm or the bSB, dSB and D-Wave Neal using the $ee$-$k_{t}$ or the angled-based QUBOs.